OpenStudy (anonymous):

find Intgrl 1/((sinx)3+(cosx)3)

OpenStudy (jwt625):

Using MATLAB: syms x;int(1/(3*sin(x)+3*cos(x))) ans = 1/3*2^(1/2)*atanh(1/4*(2*tan(1/2*x)-2)*2^(1/2))

OpenStudy (jwt625):

But if you mean sin(x)^3 and cos(x)^3 syms x;int(1/(sin(x)^3+cos(x)^3)) ans = 1/3*i*log(tan(1/2*x)^2+(1-i)*tan(1/2*x)+i)-1/3*i*log(tan(1/2*x)^2+(1+i)*tan(1/2*x)-i)+2/3*2^(1/2)*atanh(1/4*(2*tan(1/2*x)-2)*2^(1/2))